From a thin piece of cardboard 10 in. by 10 in., square corners are cut out so that the sides can be folded up to make a box. What dimensions will yield a box of maximum volume? What is the maximum volume? Round to the nearest tenth, if necessary.




A. 6.7 in. x 6.7 in. x 3.3 in.; 148.1 in³




B. 3.3 in. x 3.3 in. x 3.3 in.; 37 in³




C. 5 in. x 5 in. x 2.5 in.; 62.5 in³




D. 6.7 in. x 6.7 in. x 1.7 in.; 74.1 in³

1
Expert's answer
2021-11-02T18:29:17-0400

Let "x=" the side of cutted square corner. Then a volume of a box will e


"V=V(x)=(10-2x)^2x, \ 0\leq x\leq5"

Find the first derivative with respect to "x"


"V'(x)=((10-2x)^2x)'"

"=(10-2x)^2+2(10-2x)(-2)x"

"=(10-2x)(10-2x-4x)=4(5-x)(5-3x)"

Find the critical number(s)


"V'(x)=0=>4(5-x)(5-3x)=0"

Find the critical number(s)


"x_1=\dfrac{5}{3}, x_2=5"

"V(0)=0"

"V(5)=0"


"V(\dfrac{5}{3})=(10-2(\dfrac{5}{3}))^2(\dfrac{5}{3})=\dfrac{2000}{27}\approx74.1"

The function "V(x)" fas the absolute maximum on "[0, 5]" with value of "\dfrac{2000}{27}" at "x=\dfrac{5}{3}."


"\dfrac{5}{3}\approx1.7"


"10-2\cdot\dfrac{5}{3}=\dfrac{20}{3}\approx6.7(in)"

D. 6.7 in. x 6.7 in. x 1.7 in.; 74.1 in³